# MGAug: Multimodal Geometric Augmentation in Latent Spaces of Image Deformations

### v2

@@ IV Experimental Evaluation
IV-A Datasets
2D silhouette shapes. We choose Kimia216 [51], which is a relatively smaller dataset with seven classes of 2D silhouette shape data (bone, bottle, children, face, fish, heart, and key). Each shape class includes approximately $20$ binary images, with the size of $28^{2}$. The total number of our training images for multi-classification is about $140$.
2D handwritten digits. We use $70000$ handwritten digits of $10$ classes (00-$9$) from MNIST repository [35]. Each class includes approximately $7000$ images with the size of $28^{2}$.
3D brain MRIs. We include $803$ subjects of public T1-weighted longitudinal brain MRIs with manually delineated anatomical structures from Open Access Series of Imaging Studies (OASIS) [34], Alzheimer’s Disease Neuroimaging Initiative (ADNI) [27], Autism Brain Imaging Data Exchange (ABIDE) [14], and LONI Probabilistic Brain Atlas Individual Subject Data (LPBA40) [52]. Due to the difficulty of preserving the diffeomorphic property across individual subjects with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. All MRIs were pre-processed as $256\times 256\times 256$, $1.25mm^{3}$ isotropic voxels, and underwent skull-stripping, intensity normalization, bias field correction and pre-alignment with affine transformations.
@@ IV-B Experiments and Implementation Details
We validate the proposed model, MGAug, on both image classification and segmentation tasks. As MGAug falls into the category of transformation-based DA schemes, we mainly focuses on comparing its performance with similar baselines: randomly generated affine transformations (ATA, a commonly used augmentation strategy), and two state-of-the-art unimodal generative models based on diffeomorphic registrations (AdvEigAug or AEA, an augmentation model based on the statistics of a learned latent space of velocity fields [48]; and UDA, a unimodal deformation-based augmentation using random stationary velocity fields [13]).
@@
Evaluation of classification on small datasets. We first validate our model and other baselines on a relatively small dataset of 7-class 2D silhouette shapes. We adopt six classification network backbones, including a three-layer CNN, ResNet18 [21], VGGNet16 [54], SqueezeNet [26], DenseNet121 [25], and Vision Transformer (ViT) [17].
@@
We carefully determine the optimal number of modes, $C$, for our multimodal generative model by performing cross-validation with various values of $C$. The value of C will vary for different image analysis tasks depending on the dataset distributions. For classification, we assume that the value of C is close to the number of classes. Furthermore, we conduct a comprehensive assessment of the augmentation efficiency offered by our proposed approach. This involves augmenting the training dataset to various magnitudes, ranging from $2$ to $10$ times the size of the original dataset.
@@
Evaluation of segmentation on real 3D brain MRIs. We validate the performance of all methods on real 3D brain images with a decent size of training data. In this experiment, we employ TransUNet as our network backbone, given its proven superiority in brain segmentation performance [8]. To ensure a fair comparison of the final segmentation accuracy, we augment the training data three times for all methods, resulting in a total number of approximately $2000$ augmented images. We then compare the segmentation dice scores of all augmentation methods over different brain structures.
Evaluation of the quality of augmented images. To further evaluate the quality of the augmented images, we validate MGAug on two relatively large datasets (2D MNIST and 3D brain MRIs) with various levels of reduced training sample sizes (i.e., $10\%$, $20\%$, and $50\%$ of the original size). We augment the training images three times of the original size for all methods.
Parameter setting. We set parameter $\alpha=3.0$ for the operator $\mathcal{L}$ and the number of time steps for Euler integration in EPDiff (Eq. (3)) as $10$. We set the noise variance $\sigma$ as $0.02$, the batch size as $16$, and use the cosine annealing learning rate scheduler that starts from a learning rate $\eta=1e^{-3}$ for network training. We run all the models on training and validation images with Adam optimizer until obtaining the best validation performance. The training and prediction procedure of all learning-based methods are performed on a 40GB NVIDIA A100 Tensor Core GPUs. For both 2D and 3D datasets, we split the images by using $70\%$ as training, $15\%$ as validation, and $15\%$ as testing images.
Our method is flexible to either automatically estimated or pre-selected template(s) for group-wise data. Therefore, we demonstrate estimated with-in-class template(s) for classification tasks vs. pre-selected templates for segmentation tasks, respectively. In all experiments presented in this paper, we run cross-validation to determine the optimal number of modes $C$.
IV-C Results
@@
Tab. I reports the multi-class classification accuracy on a relatively smaller 2D silhouette shape data across all augmentation methods. For a fair comparison, we augment the data 10×10$\times$10 × of the original training images. MGAug achieves the highest performance, outperforming all baselines by $2$-$9\%$ of accuracy on six different network backbones.
The left panel of Fig. 3 reports our model MGAug’s classification accuracy (using ViT as the network backbone) on 2D silhouette shape dataset with different number of modes, $C$, along with diverse levels of augmented images quantities (i.e., 5×,7×,5$\times$,7$\times$,5 × , 7 × , and 10×10$\times$10 × the original training data size). It suggests that the utilization of a multimodal latent space $(C>1)$ consistently yields superior performance compared to an unimodal distribution $(C=1)$ by more than $2$-$7\%$. Our model achieves an optimal performance when $C=7$, which represents the number of classes in the training data.
The right panel of Fig. 3 shows a comprehensive evaluation of classification accuracy across all augmentation methods with an increased number of augmented images in training data. Our method significantly outperforms other baselines by more than $7\%$ of accuracy with different numbers of augmented images. Another interesting observation is that our MGAug achieves comparable classification accuracy when using roughly 3×3$\times$3 × augmented images for training vs. other baselines using 10×10$\times$10 × of ground truth images. This indicates the potency of our method in efficiently leveraging augmented data to achieve competitive results.
@@
Figure 3: Left: Accuracy comparison on various modes $C$ on 2D shape dataset. Right: A comparison of classification performance for all models over increasing number of augmented data taking all ground truth images.
@@
Figure 4: Examples of DetJac generated by MGAug. Left to right: source/template, target images, augmented DetJac overlaid with template images.
Fig. 5 reports the statistics of dice scores (mean and variance) of all augmentation methods over fifteen brain structures experimenting on $30\%$ of training images with $2000$ augmented images. Our MGAug outperforms all the baselines with a significant dice improvement in all anatomical structures (at least $2\%$ or higher).
Figure 5: Dice score comparison over fifteen anatomical brain structures, (LWM - Left White Matter, RWM - Right White Matter, LLV - Left Lateral Ventricle, RLV - Right Lateral Ventricle, LCC - Left Cerebellum Cortex, RCC - Right Cerebellum Cortex, LT - Left Thalamus, RT - Right Thalamus, LC - Left Cudate, RC - Right Caudate, LP - Left Putamen, RP - Right Putamen, LH - Left Hippocampus, RH - Right Hippocampus, BS - Brain Stem) for four augmentation methods.
The left panel of Fig. 6 showcases the classification results on the 2D MNIST dataset for all methods, using different percentages of ground truth data during the training phase. Our MGAug method consistently achieves the highest accuracy across all augmentation techniques. Remarkably, MGAug reaches an accuracy of $98.02\%$ (which is the baseline accuracy without any augmentation) using only $20\%$ of the ground truth data, whereas other methods require at least $50\%$ of ground truth data to achieve similar results. Similarly, the right panel of Fig. 6 presents the dice evaluation for 3D brain MRI segmentations, as the percentage of ground truth data used in training is progressively increased. In a manner mirroring the classification results, our MGAug consistently outperforms the baseline algorithms by yielding the highest dice scores. These results indicate that our augmentation method has great potential to better alleviate the heavy need for labor and time-consuming ground-truth labels.
@@
Figure 6: Left to right: classification evaluation on 2D MNIST dataset; segmentation evaluation on 3D brain MRIs with various percentages of ground truth used in the training inference. The horizontal dashed line (- -) serves as a reference, representing the classification accuracy or dice score achieved when utilizing all available training images without any augmentation.
Additional discussion.
It is to our attention that diffusion models have recently become a powerful generative model [24]. In this context, we initiated a preliminary comparison between our MGAug model and data augmentation through a diffusion model operating in the image space (by manipulating image intensities and textures) [56]. This comparison was carried out with a specific focus on classification tasks using the 2D MNIST dataset.
Experimental results show that the diffusion model achieves an accuracy of $97.56\%$ when using $50\%$ of the ground truth labels. This accuracy, interestingly, is slightly lower than what our MGAug method accomplishes ($98.02\%$) when only using $20\%$ of the ground truth data. This may indicate the advantages of multimodal distributions of latent geometric space developed in our model. Our future work will further explore the power of diffusion models in the latent geometric space for efficient data augmentation, which has not been investigated so far.
@@
MGAug vs. intensity-based augmentations. While the primary emphasis of this paper does not fall into the real of intensity-based augmentations, we compare MGAug with a few notable DA schemes manipulationg image intensities or textures. These methods include AutoAugment, an automated scheme that searches for improved augmentation policies [11]; RandAugment, an augmentation algorithm which reduces the computational expense by a simplified search space [12]; AugMix, a data processing method to improve the robustness and uncertainty estimates [22]; and TrivialAugment, a parameter-free data augmentation model [44].
@@
Tab. II reports the accuracy comparison of a variety of DA schemes for image classification networks based on the ViT backbone. The testing results on both MNIST digits and 2D shape data show that our model outperforms the baselines consistently.
@@
TABLE II: Accuracy comparison on both 2D digits and 2D shape datasets over a mix of augmentation methods (including intensity-based vs. transformation-based DA) on ViT backbone.
@@
V Conclusion
In this paper, we present a novel data augmentation model, MGAug, which learns augmenting transformations in a multimodal latent space of geometric deformations reparameterized by low-dimensional tangent spaces of diffeomorphisms. In contrast to existing augmentation models that adopt a unimodal distribution to augment data, our augmentation method fully characterizes the multimodal distribution of image transformations via introducing a mixture of multivariate Gaussians as the prior. Our method significantly improves the performance of multi-class classification (on 2D datasets) and image segmentation (on 3D brain MRIs) and eliminates the artifacts of geometric transformations by preserving the topological structures of augmented images. While MGAug focuses on LDDMM, the theoretical tool developed in our model is widely applicable to various registration frameworks, for example, stationary velocity fields [41, 19].

### v3

@@ IV-A Datasets
2D silhouette shapes
We choose Kimia216 [56], which is a relatively smaller dataset with seven classes of 2D silhouette shape data (bone, bottle, children, face, fish, heart, and key). Each shape class includes approximately $20$ binary images, with the size of $28^{2}$. The total number of our training images for multi-classification is about $140$.
2D handwritten digits
We use $70000$ handwritten digits of $10$ classes (00-$9$) from MNIST repository [40]. Each class includes approximately $7000$ images with the size of $28^{2}$.
3D brain MRIs
We include $3338$ publicly available T1-weighted 3D brain MRIs from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) [31].All subjects ranged in age from $50$ to $100$, covering cognitively normals (CN), patients affected by Alzheimer’s disease (AD), individuals with mild cognitive impairment MCI, respectively. All MRIs were preprocessed to be the size of $128\times 128\times 128$, and underwent skull-stripping, intensity min-max normalization, bias-field correction, and affine registration [54].
3D brain MRIs w/ segmentations
We include $803$ subjects of public T1-weighted longitudinal brain MRIs with manually delineated anatomical structures from Open Access Series of Imaging Studies (OASIS) [39], Alzheimer’s Disease Neuroimaging Initiative (ADNI) [31], Autism Brain Imaging Data Exchange (ABIDE) [15], and LONI Probabilistic Brain Atlas Individual Subject Data (LPBA40) [57]. Due to the difficulty of preserving the diffeomorphic property across individual subjects with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. All MRIs were pre-processed as $256\times 256\times 256$, $1.25mm^{3}$ isotropic voxels, and underwent skull-stripping, intensity normalization, bias field correction and pre-alignment with affine transformations.
@@ IV-B Experiments and Implementation Details
We validate the proposed model, MGAug, on both image classification and segmentation tasks. As MGAug falls into the category of transformation-based DA schemes, we mainly focus on comparing its performance with similar baselines: randomly generated affine transformations (ATA, a commonly used augmentation strategy), and two state-of-the-art unimodal generative models based on diffeomorphic registrations (AdvEigAug or AEA, an augmentation model based on the statistics of a learned latent space of velocity fields [51]; and UDA, a unimodal deformation-based augmentation using random stationary velocity fields [14]).
Evaluation of classification on 2D datasets
@@
We first validate our model and other baselines on a relatively small dataset of 7-class 2D silhouette shapes. We adopt six classification network backbones, including a three-layer CNN, ResNet18 [23], VGGNet16 [59], SqueezeNet [30], DenseNet121 [29], and Vision Transformer (ViT) [18].
@@
We carefully determine the optimal number of modes, $C$, for our multimodal generative model by performing cross-validation with various values of $C$. The value of C will vary for different image analysis tasks depending on the dataset distributions. For classification, we assume that the value of C is close to the number of classes. Furthermore, we conduct a comprehensive assessment of the augmentation efficiency offered by our proposed approach. This involves augmenting the training dataset to various magnitudes, ranging from $2$ to $10$ times the size of the original dataset. We further reduce the number of ground truth images to $75\%$ and $50\%$ to experiment with the effectiveness of the augmented data on this multi-class classification task.
We further validate the effectiveness of our augmentation model by experimenting on a relatively larger handwritten mnist dataset. As classification on mnist using all $75000$ ground truths almost achieves perfect classification accuracy, we experiment with $10\%$ ground truths. First, we perform classification under MLP ($4$ linear layers) and $2$-block CNN over an increased number of augmented data from 1×1$\times$1 × to 5×5$\times$5 ×. Each CNN block is built with one $3\times 3$ convolutional layer, one $2\times 2$ pooling layer, and two linear layers with ReLU activations. Second, we further validate our augmentation model by comparing its performance across a wider range of network backbones, including ResNet18/50 [23], VGGNet16/19 [59], DenseNet121/169 [29].
Evaluation of classification on real 3D brain MRIs
We evaluate the efficiency of MGAug on real-world 3D brain images, where we identify whether the subject is healthy, diagnosed with Alzheimer’s, or early/late mild cognitive impairment. Similar to the experiential design of 2D data, we conduct comprehensive experiments of the augmentation efficiency of MGAug, augmenting the training dataset to various magnitudes (1×1$\times$1 × to 5×5$\times$5 ×) over different ground-truth settings ($25\%$, $50\%$, and $75\%$). Additionally, we examine the network performance on various numbers of modes $C$.
Evaluation of segmentation on real 3D brain MRIs
We validate the performance of all methods on real 3D brain images with a decent size of training data. In this experiment, we employ UNet [55], UNetR [21], TransUNet [10] as our network backbone, given its proven superiority in brain segmentation performance. Our goal is to demonstrate that our proposed multimodal distribution learning of the deformable geometric latent space will improve the segmentation performance, regardless of what particular network backbone is. To ensure a fair comparison of the final segmentation accuracy, we augment the training data three times for all methods, resulting in a total number of approximately $2000$ augmented images. We then compare the segmentation dice scores of all augmentation methods over different brain structures.
Evaluation of the quality of augmented images
To further evaluate the quality of the augmented images, we validate MGAug on two relatively large datasets (2D MNIST and 3D brain MRIs) with various levels of reduced training sample sizes (i.e., $10\%$, $20\%$, and $50\%$ of the original size). We augment the training images three times of the original size for all methods.
Comparison with random/adversarial augmentation
We further compare our augmentation model with two additional deformable augmentations: random deformable augmentations and adversarial data augmentation [9]. The purpose of this comparison is to highlight the effectiveness of MGAug in generating plausible and structurally consistent deformations while maintaining superior performance in preserving the integrity of medical images and improving downstream tasks, including classification and segmentations.
Statistical evaluation
We perform paired t-tests [60] and report the corresponding P-values for all models across both classification and segmentation tasks on real 3D brain MRI datasets. This analysis provides insights on whether MGAug performs significantly better or differently compared to baselines.
Parameter setting
We set parameter $\alpha=3.0$ for the operator $\mathcal{L}$ and the number of time steps for Euler integration in EPDiff (Eq. (3)) as $10$. We set the noise variance $\sigma$ as $0.02$, the batch size as $16$, and use the cosine annealing learning rate scheduler that starts from a learning rate $\eta=1e^{-3}$ for network training. We run all the models on training and validation images with Adam optimizer until obtaining the best validation performance. The training and prediction procedure of all learning-based methods are performed on a 40GB NVIDIA A100 Tensor Core GPU. For both 2D and 3D datasets, we split the images by using $70\%$ as training, $15\%$ as validation, and $15\%$ as testing images.
Template selection
Our method is flexible to either automatically estimated or pre-selected template(s) for group-wise data. For the classification tasks involving 2D shapes/images and 3D brain MRI datasets, we pre-selected class-specific templates that are most representative, aligning the number of templates with the number of classes. This choice is motivated by the principle that augmented deformations should preserve group-wise structural properties, ensuring that class-specific templates maintain the topological integrity necessary for accurate classification. For segmentation tasks, we pre-selected three templates representing the healthy, diseased, and mild cognitive impairment classes. Please note that automatically estimating template images could be an alternative in our model.
In all experiments presented in this paper, we run cross-validation to determine the optimal number of modes $C$.
V Results
@@
Tab. I reports the multi-class classification accuracy on a relatively smaller 2D silhouette shape data across all augmentation methods. For a fair comparison, we augment the data 10×10$\times$10 × of the original training images. MGAug achieves the highest performance, outperforming all baselines by $2$-$9\%$ of accuracy on six different network backbones. This further proves that our augmented images help classifiers of any size (small: CNN to large: ViT) to achieve optimal performance.
@@
Figure 3: Left: Accuracy comparison on various modes $C$ on 2D shape dataset. Right: A comparison of classification performance for all models over increasing number of augmented data taking all (top-right), $75\%$ (bottom-left), and $50\%$ (bottom-right) ground truth images.
As ViT provides the best performance compared to the other network backbones, we take this feature extractor and perform a thorough study by comparing MGAug with all baselines on an increasing number of augmented data. The left panel of Fig. 3 reports our model MGAug’s classification accuracy (using ViT as the network backbone) on 2D silhouette shape dataset with different number of modes, $C$, along with diverse levels of augmented images quantities (i.e., 5×,7×,5$\times$,7$\times$,5 × , 7 × , and 10×10$\times$10 × the original training data size). It suggests that the utilization of a multimodal latent space $(C>1)$ consistently yields superior performance compared to an unimodal distribution $(C=1)$ by more than $2$-$7\%$. Our model achieves an optimal performance when $C=7$, which represents the number of classes in the training data. The right panel of Fig. 3 shows a comprehensive evaluation of classification accuracy across all augmentation methods with an increased number of augmented images in training data. Our method significantly outperforms other baselines by more than $7\%$ of accuracy with different numbers of augmented images. Another interesting observation is that our MGAug achieves comparable classification accuracy when using roughly 3×3$\times$3 × augmented images for training vs. other baselines using 10×10$\times$10 × of ground truth images. This indicates the potency of our method in efficiently leveraging augmented data to achieve competitive results.
To further validate the effectiveness of our model on varying sizes of ground truth, we take $75\%$ and $50\%$ of ground truth images from the 2D silhouette shape dataset and perform classification under the ViT backbone. Fig. 3 depicts a classification accuracy comparison of all the augmentation models on this setting over increasing number of augmented images. Our model still achieves best results in varying number of augmented images. This further proves the efficiency of using our model when we have incomplete/missing/limited ground truth images.
@@
Figure 4: Examples of DetJac generated by MGAug. Left to right: template, augmented (deformed), DetJac overlaid with augmented images.
(a) Cross validation on $C$
(b) Varying GT images.
(c) Backbone: MLP
(d) Backbone: CNN
@@
Figure 5: Top: Accuracy comparison on various modes $C$ on 2D handwritten dataset taking $20\%$ ground-truth images over 3×3$\times$3 × augmentations (left panel) and classification evaluation over increasing ground-truth images taking 3×3$\times$3 × augmentations (right panel). Bottom: A comparison of classification performance for all models on 2D handwritten digits over an increasing amount of augmented data taking $10\%$ ground truth images under MLP (left panel) and CNN backbone (right panel). The horizontal dashed line (- -) serves as a reference, representing the classification accuracy achieved when utilizing all available training images without any augmentation.
Fig. 5 reports the ablation and classification performances on the 2D handwritten MNIST dataset. Fig. 5(a) reports the classification accuracy over an increasing number of modes, $C$. Similarly, multimodal latent space $(C>1)$ yields comparatively better performance than the unimodal distribution $C=1$ and achieves optimal accuracy when $C=10$, which represents the number of classes in the MNIST dataset. Fig. 5(b) showcases the classification results on the 2D MNIST dataset for all methods, using different percentages of ground truth data during the training phase. Our MGAug method consistently achieves the highest accuracy across all augmentation techniques. Remarkably, MGAug reaches an accuracy of $98.02\%$ (which is the baseline accuracy without any augmentation) using only $20\%$ of the ground truth data, whereas other methods require at least $50\%$ of ground truth data to achieve similar results. The bottom panel of Fig. 5 reports the accuracies on the 2D MNIST dataset using $10\%$ ground truth images over an increasing number of augmented data (1×1$\times$1 × - 5×5$\times$5 ×) under two different network backbones - MLP (Fig. 5(c)) and 2-block CNN (Fig. 5(d)). It shows that MGAug consistently outperforms all the baselines over any amount of augmented data settings. Moreover, MGAug reaches the classification accuracy which uses all ground truth images (without no augmentations) by using only 2×2$\times$2 × and 2.5×2.5$\times$2.5 × of the $10\%$ ground truth images whereas the closest unimodal model UDA requires more than 3×3$\times$3 × augmented data under both MLP and CNN backbone, respectively.
Tab. II reports the classification accuracy of unimodal UDA and multimodal MGAug under different network backbones on $2\%$ ground truth images. For this experiment, we present the results for 2×2$\times$2 × and 3×3$\times$3 × augmentations. We run the experiment three times and report the average accuracy and their variances. MGAug achieves the best performance under all network backbones with very little variance, whereas unimodal UDA has comparable accuracy but with higher variances. This further shows the generalizability power of our model MGAug under various feature extractor backbones.
TABLE II: Accuracy comparison on 2D handwritten digits over unimodal UDA model across various network backbones using $2\%$ ground truth images.
Augmentation 2×2$\times$2 × 3×3$\times$3 ×
Backbones UDA MGAug UDA MGAug
ResNet18 $97.42\pm.19$ $97.90\pm.06$ $97.89\pm.03$ $\mathbf{98.10\pm.01}$
ResNet34 $97.27\pm.30$ $93.51\pm.03$ $94.52\pm.11$ $\mathbf{95.48\pm.02}$
VGGNet16 $98.37\pm.19$ $98.58\pm.08$ $98.42\pm.06$ $\mathbf{98.64\pm.05}$
VGGNet19 $98.29\pm.30$ $98.38\pm.04$ $98.57\pm.04$ $\mathbf{98.67\pm.11}$
DenseNet121 $98.27\pm.04$ $98.48\pm.01$ $98.47\pm.03$ $\mathbf{98.60\pm.02}$
DenseNet169 $98.40\pm.07$ 98.55±.002plus-or-minus98.55.00298.$5$002 $98.67\pm.11$ $\mathbf{98.78\pm.03}$
Fig. 6 (top-panel) visualizes the comparative analysis of data augmentation methods on 3D brain MRI classification. The top row presents classification accuracy trends across different proportions of ground truth data: $25\%$ (top-left), $50\%$ (top-middle), and $75\%$ (top-right) over an increased number of augmented data (0×0$\times$0 × - 5×5$\times$5 ×) for all methods. Our model consistently demonstrates superior performance across all settings, with particularly notable improvements in scenarios with limited ground truth data ($25\%$).
Fig. 6 (bottom-panel) illustrates an extensive ablation study on the impact of hyperparameter $C$ (ranging from $1$ to $5$) on model accuracy using different augmentation ratios (1×1$\times$1 × to 5×5$\times$5 ×) on the 3D brain MRI dataset taking $75\%$ ground-truth images. This ablation study reveals the sensitivity of the model to both the amount of augmented data and the choice of $C$, with optimal performance typically achieved using moderate $C$ values ($C=2$-$C=4$) and higher augmentation ratios (3×3$\times$3 ×-5×5$\times$5 ×). The results suggest that while increasing the amount of augmented data generally improves performance, the selection of an appropriate $C$ value is crucial for maximizing the benefits of the augmentation strategy in these classification tasks.
Figure 6: Top: A comparison of classification performance for all models over increasing number of augmented data taking $25\%$ (top-left), $50\%$ (top-middle), and $75\%$ (top-right) ground truth images. Bottom: Accuracy comparison on various modes C on 3D brain MRI dataset for classification tasks.
Fig. 7 presents a comprehensive evaluation of segmentation performance across different network architectures across all data augmentations on 3D brain MRI datasets along with a comprehensive ablation study on modes $C$. The top-left subplot demonstrates the ablation study of MGAug across different modes $(C=1-4)$ on three backbone architectures (UNet, UNetR, and TransUNet), where mode $C=2/4$ consistently achieves superior performance, particularly notable in UNet and UNetR backbones. The remaining three subplots showcase segmentation performance comparisons between MGAug and other baseline augmentation methods under limited ground truth scenarios $(10-30\%)$, taking 3×3$\times$3 × augmentations of the ground truth data for fair comparison. The results demonstrate that MGAug achieves comparable or superior performance to this full-data baseline using only $20-25\%$ of ground truth data. For UNet (top-right panel), MGAug maintains a clear advantage over competing methods, achieving approximately $2\%$ higher Dice scores. In UNetR (bottom-left), MGAug rapidly reaches the full-data baseline performance with just $20\%$ of ground truth data. On TransUNet (bottom-right), MGAug demonstrates consistent superiority, outperforming other methods by up to $3\%$ in Dice score. This indicates MGAug’s exceptional efficiency in leveraging limited ground truth data through augmentation, making it particularly valuable for real-world scenarios where labeled data is scarce.
Figure 7: Top-Left: Dice comparison on various modes $C$ on 3D brain MRI dataset. Right: A comparison of segmentation performance for all models over increasing number of ground-truth data ($10\%-30\%$) under UNet (Top-Right), UNetR (Bottom-Left), and TransUNet (Bottom-Right) backbones. For fair comparison, we augment the data 3×3$\times$3 × to the ground truth images. The horizontal dashed line (- -) serves as a reference, representing the segmentation dice score achieved when utilizing all available training images without any augmentation.
Fig. 8 reports the statistics of Dice scores (mean and variance) of all augmentation methods over fifteen anatomical brain structures using UNet (top), UNetR (middle), and TransUNet (bottom) backbones. For fair comparison, we use $10\%/20\%/30\%$ ground-truth images for UNet/UNetR/TransUNet backbones, respectively. Our MGAug consistently outperforms baseline methods (ATA, UDA, AEA) across most anatomical structures, particularly showing notable improvements in challenging regions like Left/Right Caudate (LC/RC) and Left/Right Putamen (LP/RP). The box plots further demonstrate MGAug’s superior performance with generally higher mean segmentation dice and comparable or lower variance across structures, though some outliers are observed in certain regions.
Figure 8: Dice score comparison over fifteen anatomical brain structures (LWM - Left White Matter, RWM - Right White Matter, LLV - Left Lateral Ventricle, RLV - Right Lateral Ventricle, LCC - Left Cerebellum Cortex, RCC - Right Cerebellum Cortex, LT - Left Thalamus, RT - Right Thalamus, LC - Left Cudate, RC - Right Caudate, LP - Left Putamen, RP - Right Putamen, LH - Left Hippocampus, RH - Right Hippocampus, BS - Brain Stem) for four augmentation methods on 3×3$\times$3 × augmentations under UNet (top), UNetR (middle), and TransUNet (bottom) backbone. For fair comparison, we use $10\%/20\%/30\%$ ground-truth images for UNet/UNetR/TransUNet backbones, respectively.
Fig. 9 presents brain MRI segmentation results across axial, coronal, and sagittal views, comparing manually delineated ground truth with predictions from all models. It shows that MGAug achieves superior performance than baselines.
Figure 9: Left to Right: Examples of segmentation (overlaid with input images) comparisons between the ground truth (GT) and predictions of all baselines including MGAug on 3D brain MRIs over three anatomical views (Top to Bottom: Axial, Coronal, and Sagittal).
Fig. 10 visualizes segmentation comparisons of our model ($C=7$) vs. unimodal ($C=1$) across three anatomical views. We highlight small anatomical structures (outlined in redbox) that show improved performance, demonstrating the superior segmentation quality of our approach.
Figure 10: Left to Right: Segmentation comparisons between ground truth (GT), multimodal ($C=7$) and unimodal ($C=1$) MGAug with a zoomed-in view, over three anatomical views (Top to Bottom: Axial, Coronal, and Sagittal).
MGAug vs. random/adversarial deformable augmentation
Tab. III presents performance evaluations comparing three MGAug with random and adversarial augmentation models on multi-class classification and segmentation tasks using 3D real brain MRI dataset. In the classification task (Tab. III(a)), we conduct experiments with varying ground-truth (GT) percentages and augmentation ratios under the ResNet backbone. MGAug consistently outperforms both Random and AdvChain approaches across all settings, demonstrating superior performance with higher GT percentages and increased augmentation ratios. Similarly, in the segmentation task (Tab. III(b)) under UNet backbone with varying GT images with 3×3$\times$3 × augmentations, MGAug demonstrates consistently better performance across different GT, achieving higher dice scores compared to the baseline methods. These results indicate MGAug’s effectiveness in both classification and segmentation tasks, particularly in scenarios with limited ground-truth data.
TABLE III: Performance comparison of random deformable augmentation (Random), adversarial deformable augmentation (AdvChain), and MGAug on (a) multi-class classification and (b) segmentation dice on 3D brain MRI dataset.
(a) Classification accuracy comparison for $50\%$ and $75\%$ ground-truth (GT) images with an increasing number of augmentations (1×1$\times$1 × to 5×5$\times$5 ×) under the ResNet backbone. GT: Ground-truth images.
GT Augmentations Random AdvChain MGAug
$50\%$ 1×1$\times$1 × $71.76$ $73.25$ $\mathbf{74.09}$
2×2$\times$2 × $72.76$ $74.75$ $\mathbf{75.42}$
3×3$\times$3 × $73.10$ $75.08$ $\mathbf{76.08}$
4×4$\times$4 × $74.09$ $75.25$ $\mathbf{76.74}$
5×5$\times$5 × $75.08$ $76.24$ $\mathbf{77.74}$
$75\%$ 1×1$\times$1 × $74.42$ $75.75$ $\mathbf{76.08}$
2×2$\times$2 × $75.08$ $76.08$ $\mathbf{77.56}$
3×3$\times$3 × $76.24$ $77.08$ $\mathbf{77.41}$
4×4$\times$4 × $76.41$ $77.23$ $\mathbf{78.41}$
5×5$\times$5 × $77.74$ $79.07$ $\mathbf{80.86}$
(b) Segmentation dice score comparison for increasing ground-truth (GT) images ($10\%$ to $30\%$) with 3×3$\times$3 × augmentations under the UNet backbone. GT: Ground-truth images
GT Random AdvChain MGAug
$10\%$ $82.65$ $84.91$ $\mathbf{86.30}$
$20\%$ $84.13$ $86.97$ $\mathbf{88.25}$
$30\%$ $88.25$ $88.44$ $\mathbf{89.96}$
MGAug vs. intensity-based augmentations
@@
While the primary emphasis of this paper does not fall into the realm of intensity-based augmentations, we compare MGAug with a few notable DA schemes manipulating image intensities or textures. These methods include AutoAugment, an automated scheme that searches for improved augmentation policies [11]; RandAugment, an augmentation algorithm which reduces the computational expense by a simplified search space [12]; AugMix, a data processing method to improve the robustness and uncertainty estimates [24]; and TrivialAugment, a parameter-free data augmentation model [48]. We maintain the same training protocols as the original models. For a fair comparison, we augment the data 10×10$\times$10 × times for 2D shapes (with all training samples) and 3×3$\times$3 × times for 2D digits (with $50\%$ of the training images) experiments.
@@
Tab. IV reports the accuracy comparison of a variety of DA schemes for image classification networks based on the ViT backbone. The testing results on both MNIST digits and 2D shape data show that our model outperforms the baselines consistently.
@@
TABLE IV: Accuracy comparison on both 2D digits and 2D shape datasets over a mix of augmentation methods (including intensity-based vs. transformation-based DA) on ViT backbone.
@@
Statistical tests
Tab. V presents the statistical significance of MGAug’s performance through p-values from paired t-tests against three baseline methods (ATA, AEA, UDA). The left subtable shows that MGAug achieves statistically significant improvements (p $<0.05$) across all three segmentation backbones, with particularly strong significance for Unet (p $<0.003$) and UNetR (p $<0.04$). Similarly, the right subtable demonstrates that MGAug maintains statistical significance across different ground truth percentages ($25\%$, $50\%$, $75\%$) for classification tasks, with p-values consistently below $0.05$.
Model ATA AEA UDA
Unet $0.0025$ $0.0023$ $0.0123$
UNetR $0.0076$ $0.0034$ $0.0370$
TransUNet $0.0343$ $0.0197$ $0.0537$
[…]
$25$ $0.037$ $0.003$ $0.048$
$50$ $0.027$ $0.012$ $0.044$
$75$ $0.048$ $0.030$ $0.018$
TABLE V: Statistical test of MGAug: (Top) P-values from paired t-tests comparing MGAug with baseline methods performing segmentation on 3D brain MRIs under different segmentation backbones; (Bottom) P-values for different ground truth percentages performing classification on 3D brain MRIs.
VI Conclusion
In this paper, we present a novel data augmentation model, MGAug, which learns augmenting transformations in a multimodal latent space of geometric deformations. In contrast to existing augmentation models that adopt a unimodal distribution to augment data, our augmentation method fully characterizes the multimodal distribution of image transformations via introducing a mixture of multivariate Gaussians as the prior. Our method significantly improves the performance of multi-class classification (on 2D datasets) and image segmentation (on 3D brain MRIs) and eliminates the artifacts of geometric transformations by preserving the topological structures of augmented images. While MGAug focuses on LDDMM, the theoretical tool developed in our model is widely applicable to various registration frameworks, for example, stationary velocity fields [45, 20].